# Counting points on character varieties

### v1

@@ 4. Counting points on character varieties
In this section, we count points on the character varieties considered in §1.2. We treat the case of a once-punctured surface first and mention the modifications required in the multi-punctured case at the end.
Let $G$ be a connected split reductive group over $k=\overline{\mathbb{F}_{q}}$ equipped with an untwisted Frobenius $F$. We assume $Z(G)$ is connected and $\mathrm{char}(k)$ is very good for $G$. Let $T$ be a maximal split $F$-stable torus of $G$. Let $S\in T^{F}$ be a strongly regular generic element and $C=G.S$ the corresponding conjugacy class in $G$. Let $X$ denote the character variety associated to the once-punctured Riemann surface and the conjugacy class $C$. By the discussions of the previous section, we have
@@
We now use a crucial result of Deligne and Lusztig to rewrite the above sum in a more tractable form.
@@ 4.1. A theorem of Deligne and Lusztig
The above corollary implies that only principal series representations contribute to the sum in the expression (1) for $|X^{F}|$. Note that one does not need Deligne–Lusztig theory to formulate the above theorem and corollary. However, the only proof we know of uses the full power of Deligne–Lusztig theory.
@@
Recall that the induced representations $R_{T}^{G}\,\theta$ and $R_{T}^{G}\,\theta^{\prime}$ have a common constituent if and only if $\theta$ and $\theta^{\prime}$ are $W$-conjugate, cf. [DL76, Corollary 6.3]. Thus, we can associate to a principal series representation $\chi\in\mathrm{Irr}(G^{F})$ a character $\theta\in(T^{F})^{\vee}$, well-defined up to $W$-conjugacy.
@@ 4.1.2. Interlude on Pontryagin duality
Let $\mu_{\infty,p^{\prime}}(\mathbb{C})$ denote the set of roots of unity in $\mathbb{C}^{\times}$ whose order is prime to $p$. We choose, once and for all, isomorphisms
@@
As noted in [DL76, §5], this induces an isomorphism
@@
$W_{\theta}=\{w\in W\,|\,w.\theta=\theta\}.$
The Langlands dual group $G_{\theta}:=(\check{G}_{\theta})^{\check{}}$ is a standard endoscopy group for $G$, containing $T$, with Weyl group $W(G_{\theta})=W_{\theta}$.
@@
Since $G$ is assumed to have connected centre, the constituents of $R_{T}^{G}\,\theta$ are in canonical bijection with $\mathrm{Irr}(W_{\theta})$, cf. [Kilmoyer78, Corollary 4.20]. Let $\chi_{\theta,\rho}\in R_{T}^{G}\,\theta$ be the irreducible character corresponding to $\rho\in\mathrm{Irr}(W_{\theta})$. Then, the double centraliser theorem (cf. [Etingof11, Theorem 5.18.1]) gives
@@
Consider the set of pairs $(L,\rho)$, where $L$ is a standard endoscopy group for $G$ and $\rho$ is an irreducible character of $W(L)$. Note $W$ acts on this set by conjugation. Let $\mathcal{T}(G)$ denote the set of $W$-orbits. We call the elements of $\mathcal{T}(G)$ $G$-types and denote them by $\tau=[L,\rho]$. [^6] Here, for ease of notation, we write $[L,\rho]$ for the orbit $[(L,\rho)]$. Note that $\mathcal{T}(G)$ is a finite set and depends only on the root system $\Phi$ of $G$; in particular, it is independent of the ground field.
By the above discussions, we can associate a type to every principal series representation. [^7] This is analogous to the types used in [HRV08, HLRV11, Cambo17]; see [Whitbread24, §4] for a detailed comparison between these notions. Using Lusztig’s Jordan decomposition, one can generalise the notion of types to all irreducible $G^{F}$-characters. We shall not need this more general version in the present paper. In other words, we have a map
@@ 4.1.6. Unipotent characters
As noted above, irreducible constituents of $R_{T}^{L}(1)$ are in canonical bijection with irreducible representations of $W(L)$. Let $\tilde{\rho}$ denote the unipotent (principal series) character of $L^{F}$ corresponding to $\rho\in\mathrm{Irr}(W(L))$.
@@
The relevance of $m_{\tau}(q)$ emerges in the following proposition which can be found in, e.g. [GM20, Corollary 2.6.6]:
@@
We can now rewrite the Frobenius Mass Formula (1) in terms of types. For each type $\tau\in\mathcal{T}(G)$, let $\mathrm{Irr}(G^{F})_{\tau}$ denote the set of irreducible principal series characters of type $\tau$. Then, we have
@@
Thus, computing $|{X}^{F}|$ is reduced to computing the character sums $S_{\tau}(q)$. [^8] In obtaining the above formula for $|{X}^{F}|$, we used the assumption that $S$ is strongly regular, but not the assumption that $S$ is generic.
4.2. Evaluating the character sums $S_{\tau}$
@@
Theorem 17 gives the following expression for $\chi_{\theta,\rho}$
@@
$\chi_{\theta,\rho}(S)=\dim(\rho)\sum_{w\in W/W_{\theta}}(w.\theta)(S)=\frac{\dim(\rho)}{|W_{\theta}|}\sum_{w\in W}\theta(w.S).$
@@
$S_{\tau}(q)=\frac{|W(L)||[L]|}{|W|}\sum_{\begin{subarray}{c}\theta\in\check{T}^{F}\\ W_{\theta}=W(L)\end{subarray}}\frac{\dim(\rho)}{|W_{\theta}|}\sum_{w\in W}\theta(w.S)=\frac{\dim(\rho)|[L]|}{|W|}\sum_{w\in W}\sum_{\begin{subarray}{c}\theta\in\check{T}^{F}\\ W_{\theta}=W(L)\end{subarray}}\theta(w.S).$
@@
$S_{\tau}(q)=\frac{\dim(\rho)|[L]|}{|W|}\sum_{w\in W}\alpha_{L,w.S}(q).$
Thus, the problem of computing $S_{\tau}$ reduces to evaluating the character sum $\alpha_{L,S}$. To compute the latter, we first consider an auxiliary sum $\alpha_{L,S}^{\supseteq}$ which is easier to handle.
4.3. The character sum $\alpha_{L,S}^{\supseteq}$
@@
This is a straightforward application of the fact that the Pontryagin dual of the embedding of finite abelian groups $(\check{T}^{F})^{W(L)}\hookrightarrow\check{T}^{F}$ is the canonical quotient map
$T^{F}\twoheadrightarrow\frac{T^{F}}{T^{F}\cap[L^{F},L^{F}]}.$
See [KNP23, §5.2] for details. ∎
@@
Now we use the fact $S\in[G,G]$ is assumed to be generic. This implies that $S\in[L,L]$ if and only if the root system $\Phi(L)$ is isolated in $\Phi=\Phi(G)$; i.e., $\Phi(L)$ is not contained in any proper Levi subsystem of $\Phi$. Now there are finitely many isolated standard endoscopic groups for $G$. Thus, replacing $F$ by $F^{m}$ if necessary, we may assume that $S\in[L^{F},L^{F}]$ for all isolated standard endoscopy groups. Hence, we obtain:
@@
Replacing the Frobenius $F$ by a finite power if necessary, we have:
$\alpha^{\supseteq}_{L,S}(q)=\begin{cases}|(\check{T}^{F})^{W(L)}|&\textrm{if $L$ is isolated}\\ 0&\textrm{otherwise}.\end{cases}$
@@ 4.3.2. Description of $\check{T}^{W(L)}$
It is proved in [KNP23, §5] that, under the assumption that $G$ has connected centre, we have:
@@
By Corollary 8, the prime $p=\mathrm{char}(k)$ does not divide $|\pi_{0}(Z(\check{L}))|$. Therefore, replacing $F$ by $F^{m}$ for some positive integer $m$ if necessary, we may assume that $q\equiv 1\mod|\pi_{0}(Z(\check{L}))|$. This implies that $\pi_{0}(Z(\check{L}))^{F}=\pi_{0}(Z(\check{L}))$. Thus, we obtain:
@@
Replacing $F$ by a finite power if necessary, we have
$\alpha^{\supseteq}_{L,S}(q)=\begin{cases}|\pi_{0}(Z(\check{L}))|\times|(Z(\check{L})^{\circ})^{F}|&\textrm{if $L$ is isolated}\\ 0&\textrm{otherwise}.\end{cases}$
Convention 25.
Henceforth, we assume $q$ is large enough so that the above equality holds; i.e., $q$ is large enough so that for every isolated standard endoscopy $L$ of $G$, we have $S\in[L^{F},L^{F}]$ and $q\equiv 1\mod|\pi_{0}(Z(\check{L}))|$.
@@ 4.4. Evaluating $\alpha_{L,S}$
Let $\mu$ denote the Möbius function on the (finite) poset of endoscopy subsystems of $\Phi$, ordered by inclusion. Möbius inversion implies
@@
Here, the sum is over standard endoscopy groups $L^{\prime}$ whose root system satisfies $\Phi(L^{\prime})\supseteq\Phi(L)$ and we have abused notation and written $\mu(L,L^{\prime})$ for $\mu(\Phi(L),\Phi(L^{\prime}))$.
@@
We have seen that $\alpha_{L^{\prime},S}^{\supseteq}$ is 00 unless $L^{\prime}$ is isolated, in which case
$Z(\check{L^{\prime}})^{\circ}=Z(\check{G})^{\circ}\simeq Z(G)^{\circ}=Z(G),$
since we assume that $G$ has connected centre, and
$\alpha_{L^{\prime},S}^{\supseteq}=|\pi_{0}(Z(\check{L^{\prime}}))|\times|Z(G)^{F}|.$
@@
Here, the sum runs over isolated standard endoscopy groups $L^{\prime}$ whose root systems contain $\Phi(L)$. Thus, $\nu(L)$ is an integer depending only on the root data of $L$ and $G$; in particular, it is independent of the ground field $k$ (of very good characteristic). In fact, $\nu(L)$ depends only on the $W$-orbit of $L$. In view of the above discussions, we conclude
(5) $\boxed{\,\alpha_{L,S}=|Z^{F}|\,\nu(L).\,}$
@@
4.5. Conclusion of the point count for the once-punctured case
We have seen that
$|{X}^{F}|=\frac{|Z^{F}|}{|T^{F}|}\sum_{\tau=[L,\rho]\in\mathcal{T}(G)}S_{\tau}(q)\,m_{\tau}(q)^{2g-1}.$
Moreover,
$S_{\tau}(q)=\frac{\dim(\rho)\,|[L]|}{|W|}\sum_{w\in W}\alpha_{L,w.S}(q)=\dim(\rho)\,|[L]|\,\nu(L)\,|Z^{F}|.$
Thus, we conclude that $|X^{F}|$ is a rational function in $q$ and that $|X^{F^{m}}|$ is obtained by evaluating the same rational function at $q^{m}$. It follows that this rational function is actually a polynomial. Thus, ${X}$ is polynomial count, in the sense of [LetellierVillegas], and our expression for $|{X}^{F}|$ gives the $E$-polynomial of $X$.

### v2

@@ 4. Counting points on character varieties
In this section, we count points on the character varieties introduced in §1.2 and prove Theorem 2. We treat the case of a once-punctured surface first and mention the modifications required in the multi-punctured case at the end.
Let $G$ be a connected split reductive group over $k=\overline{\mathbb{F}_{q}}$ equipped with an untwisted Frobenius $F$, giving $G$ an $\mathbb{F}_{q}$-structure. We assume $Z(G)$ is connected and $p=\mathrm{char}(k)$ is very good for $G$. Let $T$ be a maximal split $F$-stable torus of $G$. Let $S\in T^{F}$ be a strongly regular generic element and $C=G\cdot S$ the corresponding conjugacy class in $G$. Let $X$ denote the character variety associated to the once-punctured Riemann surface and the conjugacy class $C$.
By the Frobenius Mass Formula (Theorem 18), we have
@@
Note that the equality continues to hold if $F$ is replaced by $F^{m}$ for a positive integer $m$. In what follows, we use a crucial result of Deligne and Lusztig to rewrite the above sum in a more tractable form.
@@ 4.1. A theorem of Deligne and Lusztig
The above corollary implies that only principal series representations contribute to the sum in the expression (1) for $|X^{F}|$. Note that one does not need Deligne–Lusztig theory to formulate the above theorem and its corollary. However, the only proof we know of uses the full power of the Deligne–Lusztig theory.
@@
As we have assumed that $G$ has connected centre, the induced representations $R_{T}^{G}\,\theta$ and $R_{T}^{G}\,\theta^{\prime}$ have a common constituent if and only if $\theta$ and $\theta^{\prime}$ are $W$-conjugate, cf. [DL76, Corollary 6.3]. Thus, we can associate to a principal series representation $\chi\in\mathrm{Irr}(G^{F})$ a character $\theta=\theta_{\chi}\in(T^{F})^{\vee}$, well-defined up to $W$-conjugacy.
@@ 4.1.2. Interlude on Pontryagin duality
Let $\mu_{\infty,p^{\prime}}(\mathbb{C})$ denote the set of roots of unity in $\mathbb{C}^{\times}$ whose order is prime to $p$. We choose, once and for all, group isomorphisms
@@
As noted in [DL76, §5], this induces a group isomorphism
@@
$W_{\theta}=\{w\in W\,|\,w\cdot\theta=\theta\}.$
The Langlands dual group $G_{\theta}:=(\check{G}_{\theta})^{\check{}}$ is a standard endoscopy group for $G$ with Weyl group $W(G_{\theta})=W_{\theta}$.
@@
Since $G$ is assumed to have connected centre, the constituents of $R_{T}^{G}\,\theta$ are in canonical bijection with $\mathrm{Irr}(W_{\theta})$, cf. [Kilmoyer78, Corollary 4.20] or [DL76, Theorem 6.8]. Let $\chi_{\theta,\rho}\in R_{T}^{G}\,\theta$ be the irreducible character corresponding to $\rho\in\mathrm{Irr}(W_{\theta})$. Then, the double centraliser theorem (cf. [Etingof11, Theorem 5.18.1]) gives
@@
Consider the set of pairs $(L,\rho)$, where $L$ is a standard endoscopy group for $G$ and $\rho$ is an irreducible character of $W(L)$. Note $W$ acts on this set by conjugation. Let $\mathcal{T}(G)$ denote the set of $W$-orbits. We call the elements of $\mathcal{T}(G)$ $G$-types and denote them by $\tau=[L,\rho]$. Note that $\mathcal{T}(G)$ is a finite set which depends only on the root system $\Phi$ of $G$; in particular, it is independent of the ground field.
By the above discussions, we can associate a type to every principal series representation. (This is a reductive analogue of the notion of type used in [HRV08, HLRV11, Cambo17].) In other words, we have a map
@@ 4.1.6. Unipotent characters
Now let $L$ be a standard endoscopy subgroup of $G$. The irreducible constituents of $R_{T}^{L}1$ (which are known as unipotent principal series representations of $L^{F}$) are in bijection with irreducible characters of $W(L)$. Let $\tilde{\rho}$ denote the unipotent principal series character of $L^{F}$ corresponding to $\rho\in\mathrm{Irr}(W(L))$.
@@
The relevance of $m_{\tau}(q)$ emerges in the following proposition which can be found in, for instance, [GM20, Corollary 2.6.6]:
@@
Remark 26.
As noted in [GM20, Remark 2.3.7], for all irreducible characters $\chi$, we have $|T^{F}|$ divides $m_{\tau}(q)$.
@@
We can now rewrite the Frobenius Mass Formula (1) in terms of types. To this end, let $\mathrm{Irr}(G^{F})_{\tau}$ denote the set of irreducible principal series characters of type $\tau$. Then, we have
@@
The importance of the above formula is that the sum is over a parameter which does not depend on $q$. Moreover, it implies that computing $|{X}^{F}|$ is reduced to computing the character sums $S_{\tau}(q)$.
Remark 27.
  1. (i)
    In obtaining the above formula, we used the assumption that $S$ is strongly regular, but not the assumption that $S$ is generic.
  2. (ii)
    The above formula holds if $F$ is replaced by $F^{m}$ for a positive integer $m$.
4.2. Evaluating the character sum $S_{\tau}$
@@
Next, Theorem 22 implies
@@
$\chi_{\theta,\rho}(S)=\dim(\rho)\sum_{w\in W/W_{\theta}}(w\cdot\theta)(S)=\frac{\dim(\rho)}{|W_{\theta}|}\sum_{w\in W}\theta(w\cdot S).$
@@
$S_{\tau}(q)=\frac{|W(L)||[L]|}{|W|}\sum_{\begin{subarray}{c}\theta\in\check{T}^{F}\\ W_{\theta}=W(L)\end{subarray}}\frac{\dim(\rho)}{|W_{\theta}|}\sum_{w\in W}\theta(w\cdot S)=\frac{\dim(\rho)|[L]|}{|W|}\sum_{w\in W}\sum_{\begin{subarray}{c}\theta\in\check{T}^{F}\\ W_{\theta}=W(L)\end{subarray}}\theta(w\cdot S).$
@@
$S_{\tau}(q)=\frac{\dim(\rho)|[L]|}{|W|}\sum_{w\in W}\alpha_{L,w\cdot S}(q).$
This formula continues to hold if $q$ is replaced by $q^{m}$ for a positive integer $m$.
The above equality implies that the problem of computing $S_{\tau}$ reduces to evaluating the character sum $\alpha_{L,S}$. To compute the latter, we first consider an auxiliary sum $\alpha_{L,S}^{\supseteq}$ which is easier to handle.
4.3. The auxiliary sum $\alpha_{L,S}^{\supseteq}$
@@
The lemma follows from the fact (proved in [KNP23, §5.2]) that the Pontryagin dual of the embedding of finite abelian groups $(\check{T}^{F})^{W(L)}\hookrightarrow\check{T}^{F}$ is the canonical quotient map
$T^{F}\twoheadrightarrow\frac{T^{F}}{T^{F}\cap[L^{F},L^{F}]}.\qed$
@@
Now we simplify the above expression using the fact $S\in[G,G]$ is assumed to be generic. This implies that $S\in[L,L]$ if and only if the root system $\Phi(L)$ is isolated in $\Phi=\Phi(G)$; i.e., $\Phi(L)$ is not contained in any proper Levi subsystem of $\Phi$.
Note that the containment $[L^{F},L^{F}]\subseteq[L,L]$ may be proper (e.g. $L=G=\mathrm{PGL}_{n}$). Thus, it could happen that $S\in[L,L]\setminus[L^{F},L^{F}]$. However, this subtlety goes away after a finite base change. Indeed, there are finitely many isolated standard endoscopic groups for $G$; thus, given $S\in T^{F}$, there exists a positive integer $m$, such that $S\in[L^{F^{m}},L^{F^{m}}]$ for all isolated standard endoscopy groups. We conclude:
@@
There exists a positive integer $m$ such that
$\alpha^{\supseteq}_{L,S}(q^{m})=\begin{cases}|(\check{T}^{F^{m}})^{W(L)}|&\textrm{if $L$ is isolated}\\ 0&\textrm{otherwise}.\end{cases}$
Moreover, the same formula holds if $m$ is replaced by $mx$ for a positive integer $x$.
@@ 4.3.2. Description of $\check{T}^{W(L)}$
To proceed further, we need a description of $T^{W(L)}$. It is proved in [KNP23, §5] that, under the assumption that $G$ has connected centre, we have:
@@
By Corollary 11, the prime $p=\mathrm{char}(k)$ does not divide $|\pi_{0}(Z(\check{L}))|$. Thus, there exists a positive integer $n$ such that $q^{n}\equiv 1\mod|\pi_{0}(Z(\check{L}))|$. This, in turn, implies
$\pi_{0}(Z(\check{L}))^{F^{n}}=\pi_{0}(Z(\check{L})).$
Combining this with the previous corollary, we obtain:
@@
There exists a positive integer $r$ such that for every standard endoscopy $L$ of $G$, we have
$\alpha^{\supseteq}_{L,S}(q^{r})=\begin{cases}|\pi_{0}(Z(\check{L}))|\times|(Z(\check{L})^{\circ})^{F^{r}}|&\textrm{if $L$ is isolated}\\ 0&\textrm{otherwise}.\end{cases}$
Moreover, the same formula holds if $r$ is replaced by $rx$ for any positive integer $x$.
@@ 4.4. Evaluating $\alpha_{L,S}$
Let $\mu$ denote the Möbius function on the (finite) poset of the standard endoscopy groups of $G$, ordered by inclusion of their root systems (Note that these root systems all live in $\Phi(G)$). Möbius inversion implies
@@
where the sum is over standard endoscopy groups $L^{\prime}$ satisfying $\Phi(L^{\prime})\supseteq\Phi(L)$.
@@
We have seen that $\alpha_{L^{\prime},S}^{\supseteq}$ is 00 unless $L^{\prime}$ is isolated. In this case, our assumption that $G$ has connected centre implies
$Z(\check{L^{\prime}})^{\circ}=Z(\check{G})^{\circ}\simeq Z(G)^{\circ}=Z(G).$
@@
Here, the sum runs over isolated standard endoscopy groups $L^{\prime}$ of $G$ whose root systems contain $\Phi(L)$. Thus, $\nu(L)$ is an integer depending only on the root data of $L$ and $G$; in particular, it is independent of the ground field $k$ (of very good characteristic). In fact, $\nu(L)$ depends only on the $W$-orbit of $L$. In view of the above discussions, we conclude:
Proposition 32.
There exists a positive integer $r$ such that
$\alpha_{L,S}(q^{r})=|Z^{F^{r}}|\,\nu(L).$
Moreover, the equality continues to hold if $r$ is replaced by $rx$ for any positive integer $x$.
@@
4.5. Proof of Theorem 2 in the once-punctured case
The above discussions imply that there exists a positive integer $r$ such that
$|{X}(\mathbb{F}_{q^{r}})|=\frac{|Z(\mathbb{F}_{q^{r}})|}{|T(\mathbb{F}_{q^{r}})|}\sum_{\tau=[L,\rho]\in\mathcal{T}(G)}S_{\tau}(q^{r})\,m_{\tau}(q^{r})^{2g-1},$
where
$S_{\tau}(q^{r})=\frac{\dim(\rho)\,|[L]|}{|W|}\sum_{w\in W}\alpha_{L,w\cdot S}(q^{r})=\dim(\rho)\,|[L]|\,\nu(L)\,|Z(\mathbb{F}_{q^{r}})|.$
Moreover, the above formulas continue to hold if $r$ is replaced by a positive multiple.
4.5.1.
To proceed further, we need some notation. Let $f_{Z}$, $f_{T}$, $f_{L}$ be counting polynomial for $Z$, $T$, $L$. Thus,
$f_{Z}(t)=(t-1)^{\dim(Z)},\quad f_{T}(t)=(t-1)^{\dim(T)},\quad f_{L}(t)=f_{T}(t)t^{|\Phi^{+}(L)|}P_{L}(t),$
where $P_{L}(t)\in\mathbb{Z}[t]$ is the Poincaré polynomial of the flag variety of $L$. Recall that for a type $\tau=[L,\rho]$, the polynomial $m_{\tau}(t)$ is defined by
$m_{\tau}(t):=t^{|\Phi^{+}(G)|-|\Phi^{+}(L)|}\frac{|f_{L}(t)|}{\tilde{\rho}(1)}\in\mathbb{Q}[t].$
Definition 33.
Define the rational function
$f_{X}(t):=\frac{f_{Z}(t)^{2}}{f_{T}(t)}\sum_{\tau=[L,\rho]\in\mathcal{T}(G)}\dim(\rho)|[L]|\,m_{\tau}(t)^{2g-1}\in\mathbb{Q}(t).$
4.5.2. Conclusion of the proof
We have seen that there exists a positive integer $r$ such that
$|X(\mathbb{F}_{q^{rm}})|=f_{X}(q^{rm}),\qquad\forall\,m\in\mathbb{Z}_{>0}.$
Thus, $f_{X}(q^{rm})$ is an integer for all $m>0$. An elementary argument (cf. [GoodwinRohrle, Lemma 2.11]) then implies that $f_{X}(t)$ is a polynomial. Thus, $X\times_{\mathbb{F}_{q}}\mathbb{F}_{q^{r}}$ is polynomial count with counting polynomial $f_{X}$. This concludes the proof of Theorem 2 in the single punctured case.
4.6. Proof of Theorem 2 in the multi-punctured case
Let ${X}$ be a character variety associated to an $n$-punctured surface group satisfying Assumption 1. Define the rational function
(5) $f_{X}(t):=\frac{|W|^{n-1}f_{Z}(t)^{2}}{f_{T}(t)^{n}}\sum_{\tau=[L,\rho]\in\mathcal{T}(G)}\frac{\dim(\rho)^{n}|[L]|\,m_{\tau}(t)^{2g-2+n}}{|W(L)|^{n-1}}\in\mathbb{Q}(t).$
Our goal is to show that
$|X(\mathbb{F}_{q^{rm}})|=f_{X}(q^{rm}),\qquad\forall\,m\in\mathbb{Z}_{>0}.$
4.6.1.
As in the once-punctured case, we have
$|X^{F}|=\frac{|Z^{F}|}{|T^{F}|^{n}}\sum_{\tau=[L,\rho]\in\mathcal{T}(G)}m_{\tau}(q)^{2g-2+n}S_{\tau}(q),$
where $S_{\tau}$ is now defined as
$S_{\tau}(q):=\sum_{\chi\in\mathrm{Irr}(G^{F})_{\tau}}\chi(S_{1})\cdots\chi(S_{n}).$
4.6.2.
Proceeding as in the once-punctured case, we compute
$S_{\tau}(q)=\frac{|W(L)|}{|N_{W}(W(L))|}\sum_{\begin{subarray}{c}\theta\in\check{T}^{F}\\ W_{\theta}=W(L)\end{subarray}}\prod_{i=1}^{n}\chi_{\theta,\rho}(S_{i})=\frac{|W(L)|}{|W|}|[L]|\sum_{\begin{subarray}{c}\theta\in\check{T}^{F}\\ W_{\theta}=W(L)\end{subarray}}\prod_{i=1}^{n}\chi_{\theta,\rho}(S_{i}).$
We already computed
$\chi_{\theta,\rho}(S_{i})=\frac{\dim(\rho)}{|W_{\theta}|}\sum_{w\in W}\theta(w\cdot S_{i}).$
Plugging this expression for $\chi_{\theta,\rho}(S_{i})$ into the one for $S_{\tau}$ yields
$S_{\tau}(q)=\frac{\dim(\rho)^{n}\,|[L]|}{|W|\,|W(L)|^{n-1}}\sum_{\underline{w}\in W^{n}}\alpha_{L,\underline{w}\cdot\underline{S}}(q),$
where $\underline{w}:=(w_{1},\ldots,w_{n})\in W^{n}$ and $\underline{w}\cdot\underline{S}:=(w_{1}\cdot S_{1})\cdots(w_{n}\cdot S_{n})\in T^{F}$.
4.6.3.
We have seen that there exists a positive integer $r$ such that
$S_{\tau}(q^{r})=\frac{|W|^{n-1}}{|W(L)|^{n-1}}\dim(\rho)^{n}\,|[L]|\,\nu(L)\,|Z^{F^{r}}|.$
This means we have
(6) $|X^{F^{r}}|=\frac{|W|^{n-1}|Z^{F^{r}}|^{2}}{|T^{F^{r}}|^{n}}\sum_{\tau=[L,\rho]\in\mathcal{T}(G)}\frac{\dim(\rho)^{n}\,|[L]|\,m_{\tau}(q^{r})^{2g-2+n}}{|W(L)|^{n-1}}=f_{X}(q^{r}).$
The conclusion continues to hold if $r$ is replaced by a positive integer multiple $mr$. The same reasoning as in the single punctured case concludes the proof of Theorem 2. ∎
5. Topological implications
The aim of this section is discuss the topological implications of our expression for the counting polynomial of generic character varieties and prove Corollary 3. We continue using the notation of the previous section.